# Double Robustness of Local Projections

### v3

@@ A.3 Covariance structure of LP and VAR estimators
and the $\operatorname{aCov}$” notation refers to elements of the asymptotic variance-covariance matrix in LABEL:eqn:asy_normal_joint in LABEL:app:joint. In particular, $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}}-\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})=0$.
@@ Appendix B Proofs
Lemmas whose name begins with “E” can be found in LABEL:app:proof_details.
@@
LABEL:thm:lp_coefficients_aux shows that we can represent
@@
where the expressions for the coefficient matrices and the $1\times n$ two-sided lag polynomial $\Theta_{h}(L)=\sum_{\ell=-\infty}^{\infty}\Theta_{h,\ell}L^{\ell}$ are given in LABEL:thm:lp_coefficients_aux.
@@
[…]
@@
by LABEL:thm:x(LABEL:itm:thm_x_iv) and (LABEL:itm:thm_x_v)
[…]
@@
LABEL:thm:A shows that $\hat{A}-A=O_{p}(T^{-\zeta}+T^{-1/2})$. By Magnus and Neudecker (2007, Table 7, p. 208),
@@
where $\Psi_{h}\equiv\sum_{\ell=1}^{h}A^{h-\ell}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{\ell-1}$. LABEL:thm:A further implies that
@@
where $S$ was defined in Assumption˜2.1. LABEL:thm:nu shows that

### v4

@@ A.3 Covariance structure of LP and VAR estimators
and the $\operatorname{aCov}$” notation refers to elements of the asymptotic variance-covariance matrix in Equation (C.2) in Supplemental Appendix C.4. In particular, $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}}-\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})=0$.
@@ Appendix B Proofs
Lemmas whose name begins with “E” can be found in Supplemental Appendix E.
@@
Lemma E.1 shows that we can represent
@@
where the expressions for the coefficient matrices and the $1\times n$ two-sided lag polynomial $\Theta_{h}(L)=\sum_{\ell=-\infty}^{\infty}\Theta_{h,\ell}L^{\ell}$ are given in Lemma E.1.
@@
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}(y_{i^{*},t+h}-\theta_{h,T}\hat{x}_{h,t}-\underline{B}_{h,y}^{\prime}\underline{y}_{j^{*},t}-B_{h,y}^{\prime}y_{t-1})\hat{x}_{h,t}}{\sigma_{j^{*}}^{2}+o_{p}(1)}\quad\text{by Lemma E.4(v)}$
@@
by Lemma E.4(iii) and (vi)
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}\xi_{i^{*},h,t}\varepsilon_{j^{*},t}+O_{p}(T^{-2\zeta})+o_{p}(T^{-1/2})}{\sigma_{j^{*}}^{2}+o_{p}(1)}\quad\text{by Lemma E.1},$
@@
Lemma E.2 shows that $\hat{A}-A=O_{p}(T^{-\zeta}+T^{-1/2})$. By Magnus and Neudecker (2007, Table 7, p. 208),
@@
where $\Psi_{h}\equiv\sum_{\ell=1}^{h}A^{h-\ell}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{\ell-1}$. Lemma E.2 further implies that
@@
where $S$ was defined in Assumption˜2.1. Lemma E.3 shows that